An 11 m ladder is placed 4 m away from the wall. If Joey, who is 1.7 m tall, stood up on top of the ladder,

how high would the top of his head be in relation to the floor?


1
Expert's answer
2020-11-04T17:08:48-0500



"\textsf{Ladder}= 11m\\
\textsf{Floor} = 4m\\
\textsf{floor to Joey's feet } = y\\
\textsf{According to pythagoras theorem,}\\
4^2 + y^2 = 11^2\\
y^2 = 11^2 - 4^2\\
y^2 = 121 - 16\\
y^2 = 105\\
y = 10.25m"


From the diagram above,

"x = y + 1.7m\\
x = 10.25m + 1.7m\\
x = 11.95m\\"


"\therefore 11.95m" is how high the top of his head would be in relation to the floor.